OpenStudy (anonymous):

What is the inverse of f(x) = 2x − 8?

OpenStudy (ivettef365):

for the inverse you write problem as y = 2x - 8 isolate the x what do you get ??

OpenStudy (anonymous):

4

OpenStudy (ivettef365):

y+8 = 2x --- -- 2 2

OpenStudy (ivettef365):

so x = y+8/2, now change the y and the x and the inverse y=x+8/2